(n/2)-14 is less than or = to 5
Answer by Theo(13342)

the inequality is:

n/2 - 14 <= 5
add 14 to both sides to get:
n/2 <= 19
multiply both sides by 2 to get:
n <= 38

that's your solution.

confirm by replacing n with 38 in the original inequality to get:
n/2 - 14 <= 5 becomes 38/2 - 14 <= 5 which becomes 19 - 14 <= 5 which becomes 5 <= 5 which is true.

try it with n = 36 which is less than 38 to get:
n/2 - 14 <= 5 becomes 36/2 - 14 <= 5 which becomes 18 - 14 <= 5 which becomes 4 <= 5 which is true.

try it with n = 40 which is greater than 38 to get:
n/2 - 4 <-= 5 becomes 40/2 - 14 <= 5 which becomes 20 - 14 <= 5 which becomes 6 <= 5 which is false.

the inequality is only true when n is smaller than or equal to 38.

the solution is confirmed to be good.